Evaluate the following system of equation by Gaussian Elimination method

x1+5x2+2x3=9

x1+x2+7x3=6

-3x2+4x3=-2



1
Expert's answer
2021-06-03T17:01:50-0400

Augmented matrix


"A=\begin{pmatrix}
 1 & 5 & 2 & \ \ 9\\
 1 & 1 & 7 & \ \ 6\\
 0 & -3 & 4 & \ \ -2\\
\end{pmatrix}"

"R_2=R_2-R_1"


"\begin{pmatrix}
 1 & 5 & 2 & \ \ 9\\
 0 & -4 & 5 & \ \ -3\\
 0 & -3 & 4 & \ \ -2\\
\end{pmatrix}"

"R_2=-\dfrac{R_2}{4}"


"\begin{pmatrix}
 1 & 5 & 2 & \ \ 9\\
 0 & 1 & -5/4 & \ \ 3/4\\
 0 & -3 & 4 & \ \ -2\\
\end{pmatrix}"

"R_1=R_1-5R_2"


"\begin{pmatrix}
 1 & 0 & 33/4 & \ \ 21/4\\
 0 & 1 & -5/4 & \ \ 3/4\\
 0 & -3 & 4 & \ \ -2\\
\end{pmatrix}"

"R_3=R_3+3R_2"


"\begin{pmatrix}
 1 & 0 & 33/4 & \ \ 21/4\\
 0 & 1 & -5/4 & \ \ 3/4\\
 0 & 0 & 1/4 & \ \ 1/4\\
\end{pmatrix}"

"R_3=4R_3"


"\begin{pmatrix}
 1 & 0 & 33/4 & \ \ 21/4\\
 0 & 1 & -5/4 & \ \ 3/4\\
 0 & 0 & 1 & \ \ 1\\
\end{pmatrix}"

"R_1=R_1-\dfrac{33R_3}{4}"


"\begin{pmatrix}
 1 & 0 & 0 & \ \ -3\\
 0 & 1 & -5/4 & \ \ 3/4\\
 0 & 0 & 1 & \ \ 1\\
\end{pmatrix}"

"R_2=R_2+\dfrac{5R_3}{4}"


"\begin{pmatrix}
 1 & 0 & 0 & \ \ -3\\
 0 & 1 & 0 & \ \ 2\\
 0 & 0 & 1 & \ \ 1\\
\end{pmatrix}"


"x_1=-3, x_2=2, x_3=1"

"(-3,2,1)"